I need help figuring out this fraction problem 16 1/2 / 9 1/2=
Answer by Alan3354(69443)

(16 1/2) / (9 1/2)
The easy way:
Multiply NUM and DEN by 2
= 33/19
That's it.
-------------
The usual way:
= (33/2) / (19/2)
Invert the divisor and multiply
= (33/2) * (2/19)
= 66/38
= 33/19
--------
Or 1 14/19, tho a mixed number is useless.